# WKB structure in a scalar model of flat bands

### v1

@@ 5. A toy model
Under an additional assumption it is easy to that there are no protected states at $k\in\Lambda^{*}$:
@@
$\begin{gathered}\begin{pmatrix}(2D_{\bar{z}})^{2}&R_{-}\\ R_{+}&0\end{pmatrix}^{-1}:H^{1}(\mathbb{C}/\Lambda;\mathbb{C})\oplus\mathbb{C}\to L^{2}(\mathbb{C}/\Lambda;\mathbb{C})\oplus\mathbb{C},\\ R_{-}u_{-}=u_{-}\ \ \ R_{+}u=\langle u,1\rangle.\end{gathered}$
@@
$\begin{pmatrix}E(\alpha)&E_{+}(\alpha)\\ E_{-}(\alpha)&E_{-+}(\alpha)\end{pmatrix},\ \ E_{-+}(\alpha)=\sum_{k=1}^{\infty}(-1)^{k}\alpha^{2k}E_{-}V(EV)^{k-1}E_{+}.$ (5.3)
@@
Remark. Suppose that $\Gamma\subset\Lambda^{*}$ such that $0\notin\Gamma$ and for $\gamma,\gamma^{\prime}\in\Gamma$, $\gamma+\gamma^{\prime}\in\Gamma$. Suppose that
$V(z)=\sum_{\gamma\in\Gamma}a_{\gamma}e^{i\langle z,\gamma\rangle}\in C^{\infty}(\Lambda/\Gamma),\ \ \ a_{\gamma}=\mathcal{O}(|\gamma|^{-\infty}).$ (5.4)
@@
In view of the lemma it natural to ask about the set
@@
$m(\alpha,k):=\frac{1}{2\pi i}\operatorname{tr}_{L^{2}(\mathbb{C}/\Lambda)}\oint_{\alpha}P(\zeta,k)^{-1}\partial_{\zeta}P(\zeta,k)d\zeta,$ (5.6)
where the integral is over a positively oriented circle including $\alpha$ as the only possible pole of $\zeta\mapsto P(\zeta,k)^{-1}$. If $k\in\Lambda^{*}$ this assumes meromorphy of $\zeta\mapsto P(\zeta,k)^{-1}$.
In this section we will consider this question for the simplest case $\Lambda=\mathbb{Z}^{2}$ and $W(x,y)=W(x)$, $(x,y)\in\mathbb{R}^{2}$. That amounts to the study of the operator (1.13). We present a self-contained argument which indicates that even in the simplified setting (compared to that in §4 a rigorous derivation becomes quite involved.
@@ 5.1. Semiclassical formulation
for some constant $C_{0}$. We should the problem scales and we could consider $|\alpha|\gg 1$ instead of (5.7) and $h\to 0$. It is however natural in the context of WKB to use real $h\to 0$ as the asymptotic parameter.
@@
For $z_{0},z_{1}\in\mathbb{C}$, define the transition matrix $\mathcal{M}_{z_{0}\to z_{1}}:\mathbb{C}^{2}\to\mathbb{C}^{2}$ as follows: for each solution $u$ to (5.8) we have
@@ 5.2. WKB approximation
We will now study the solutions to (5.8) using the global phase function
@@ 5.2.2. WKB solutions
We next construct approximate WKB solutions, $u^{\pm}_{N}$, to the equation (5.8). Let $\gamma$ be a curve satisfying (5.15). Since $\gamma$ does not self ntersect, we can fix a simply connected open set set $\mathcal{U}_{\gamma}$ such that
@@
$\begin{split}\operatorname{Wr}(u_{N}^{+},u_{N}^{-})&=e^{b_{N}^{+}+b_{N}^{-}}\left(h\partial_{z}(\alpha\Phi-ihb_{N}^{+})-(-h\partial_{z}(\alpha\Phi-ihb_{N}^{-})\right)\\ &=\alpha e^{b_{N}^{+}+b_{N}^{-}}(2W-i\tilde{h}\partial_{z}(b_{N}^{+}-b_{N}^{-})).\end{split}$
On the other hand, (5.12), (5.22) and (5.23) give $\operatorname{Wr}(u_{N}^{+},u_{N}^{-})=\mathcal{O}(\tilde{h}^{N+1})$.
@@ 5.2.3. Approximating the transition matrix
$S_{\gamma}(\alpha,h):=\begin{pmatrix}e^{i(\Phi_{+}(\gamma(0);\alpha,h)-\Phi_{+}(\gamma(1);\alpha,h))/h}&0\\ 0&e^{i(\Phi_{-}(\gamma(0);\alpha,h)-\Phi_{-}(\gamma(1);\alpha,h))/h}\end{pmatrix}$. (5.29)
@@
$\mathcal{M}_{\gamma(1)\leftarrow\gamma(0)}=$ $\,B_{N}(\gamma(1);\alpha,h)^{-1}S_{\gamma}(\alpha,h)B_{N}(\gamma(0);\alpha,h)$ (5.30)
@@
$e^{-i\Phi_{+}(\gamma(1);\alpha,h)/h}\operatorname{Wr}_{+}(1)$ $=e^{(-i\Phi_{+}(\gamma(1);\alpha,h)+i\Phi_{+}(\gamma(0);\alpha,h))/h)}e^{-i\Phi_{+}(\gamma(0);\alpha,h)}\operatorname{Wr}_{+}(t)$
@@
$e^{-i\Phi_{-}(\gamma(1);\alpha,h)/h}\operatorname{Wr}_{-}(1)$ $=e^{(-i\Phi_{-}(\gamma(1);\alpha,h)+i\Phi_{-}(\gamma(0);\alpha,h))/h)}e^{-i\Phi_{-}(\gamma(0);\alpha,h)}\operatorname{Wr}_{+}(t)$
@@
(we used here the facts that $\operatorname{Re}(-i\Phi_{\pm}(\gamma(1);\alpha,h)=\pm\psi_{\gamma,\alpha}(1)+\mathcal{O}(h)$) that is,
${\rm{\bf{Wr}}}_{N}(\gamma(1))=S_{\gamma}(\alpha,h){\rm{\bf{Wr}}}_{N}(\gamma(0))+\mathcal{O}_{\gamma,N}(h^{N+1}e^{(\psi_{\gamma,\alpha}(1)-\psi_{\gamma,\alpha}(0))/h}|\mathbf{u}(\gamma(0))|).$
@@
$B_{N}(\gamma(1);\alpha,h)\mathbf{u}(\gamma(1))=$ $S_{\gamma}(\alpha,h)B_{N}(\gamma(0);\alpha,h)\mathbf{u}(\gamma(0))$
@@
$\mathbf{u}(\gamma(1))=$ $B_{N}(\gamma(1);\alpha,h)^{-1}S_{\gamma}(\alpha,h)B_{N}(\gamma(0);\alpha,h)\mathbf{u}(\gamma(0))$
@@ 5.3. WKB analysis in the periodic case
We also assume that $\gamma$ is a curve as in (5.15) such that
@@ 5.3.2. Stokes loops
Assume that $W$ is a 1-periodic entire function on $\mathbb{C}$. Let $\gamma:[0,1]\to\mathbb{C}$ be a smooth non-self-intersecting curve. We say that $\gamma$ is a Stokes loop for $W$, if $\gamma$ extends to a smooth non-self-intersecting curve $\widetilde{\gamma}$$\mathbb{{\mathbb{R}}}\to\mathbb{C}$ and the following conditions hold:
@@
$W(\gamma(t))\gamma^{\prime}(t)$ $>0\quad\text{for all }t\in[0,1]$ (5.40)
@@
Suppose that $\gamma$ is a Stokes loop in the sense of Definition 15. Then the the winding number condition (5.34) holds.
@@
$\gamma(t):=t+if(t),\ \ f\in C^{\infty}(\mathbb{R};\mathbb{R}).$ (5.46)
@@ 5.4. Mutliplicity in terms of the transition matrix
$\begin{gathered}P_{k}(\alpha,h):H^{2}_{k}\to L^{2}_{k},\ \ \ H^{s}_{k}:=\{w\in L^{s}_{\rm{loc}}(\mathbb{R}):w(x+1)=e^{ik/h}w(x)\},\end{gathered}$ (5.48)
$(/H_{k}^{0}=L^{2}_{k}=L^{2}$($\mathbb{R}$$\mathbb{Z}$) and then put
$m_{h}(\alpha,k)=\frac{1}{2\pi i}\operatorname{tr}_{L^{2}_{k}}\oint_{\alpha}P_{k}(\zeta,h)^{-1}\partial_{\zeta}P_{k}(\zeta,h)d\zeta,$ (5.49)
where the integral is over a positively oriented circle including $\alpha$ as the only possible pole of $\zeta\mapsto P_{k}(\zeta,h)$. Since, in the notation of (1.1), $P_{k}(\zeta,h)=h^{2}P(\zeta/h,k)$ we see that
$m_{h}(\alpha,k)=m(\alpha/h,k/h),$ (5.50)
@@
We note that, in the notation of (1.13)
$\begin{gathered}P_{k}(\alpha,h)=h^{2}\tau(k)P_{k/h}(\alpha/h)\tau(k)^{-1},\\ \tau(k)w(x):=e^{ikx/h}w(x),\ \ \ \tau(k):H^{s}(\mathbb{R}/\mathbb{Z})\to H^{s}_{k}.\end{gathered}$
Hence, the meromorphy of $\zeta\mapsto P_{k}(\zeta,h)$ is guaranteed by Lemma 9, except for $k/h\in 2\pi\mathbb{Z}$ when we can have $\ker P_{k}(\alpha,h)\neq\{0\}$ for all $\alpha$ (see the remark after Lemma 9). In that case, we put $m_{h}(\alpha,k)=\infty$.
@@
$m_{h}(\alpha,k)=\frac{1}{2\pi i}\operatorname{tr}_{\mathbb{C}^{2}}\oint_{\alpha}(\mathcal{M}_{z_{0}+1\leftarrow z_{0}}(\zeta,h)-e^{ik/h})^{-1}\partial_{\zeta}\mathcal{M}_{z_{0}+1\leftarrow z_{0}}(\zeta,h)d\zeta,$ (5.51)
@@
$\zeta\mapsto\det(\mathcal{M}_{z_{0}+1\leftarrow z_{0}}(\zeta,h)-e^{ik/h})$
at $\zeta=\alpha$.
@@
We start by observing that since $W(z+1)=W(z)$, $\mathcal{M}_{z_{0}\leftarrow z}=\mathcal{M}_{z_{0}+1\leftarrow z+1}$. Hence
@@
and the statements for a fixed $z_{0}$ are equivalent to statements for all $z$. The obstruction to the meromorphy of $P_{k}(\alpha,h)^{-1}$ is having a non-trivial solution $w(\alpha)\in H_{k}^{2}$ to $P_{k}(\alpha,h)w(\alpha)=0$ for all $\alpha$. As discussed in §5.1 $x\mapsto w(\alpha,x)$ extends to an entire function which then satisfies (5.47) and $P(\alpha,h)w=0$. Definition of $\mathcal{M}_{z+1\leftarrow z}(\alpha,h)$ in §5.1, then shows that $e^{ik/h}$ is an eigenvalue of $\mathcal{M}_{z+1\leftarrow z}$ for all $z$. That gives $\Leftarrow$ in the first statement. To see $\Rightarrow$ we note that meromorphy of $P_{k}(\alpha,h)^{-1}$ implies existence of $\alpha_{0}$ for which $\ker P_{k}(\alpha_{0},h)=\{0\}$. If $e^{ik/h}$ were an eigenvalue of $\mathcal{M}_{z_{0}+1\leftarrow z_{0}}(\alpha_{0},h)$ then, by (5.52) it would be an eigenvalue of $\mathcal{M}_{z+1\leftarrow z}(\alpha_{0},h)$ for all $z$ and we would have have an entire solution to $P(\alpha_{0},h)u=0$ satisfying (5.47). But that means that $\ker P_{k}(\alpha,h)$ is nontrivial.
@@
The related operator acting on holomorphic functions is given by $\widetilde{P}_{k}(\alpha,h)=(hD_{z}+k)^{2}-\alpha^{2}$. The corresponding transition matrix, $\mathcal{M}^{k}_{z_{1}\leftarrow z_{0}}$, is defined as in (5.10) but with (5.9) replaced by $\mathbf{u}(z)=(u(z),(hD_{z}+k)u(z))$. It is related to the original transition matrix as follows:
$\mathcal{M}^{k}_{z_{1}\leftarrow z_{0}}(\alpha,h)=e^{-ik(z_{0}-z_{1})/h}\mathcal{M}_{z_{1}\leftarrow z_{0}}(\alpha,h).$ (5.53)
@@
$D_{k}(\alpha,h):=\begin{pmatrix}hD_{z}+k&-1\\ -V&hD_{z}+k\end{pmatrix},\ \ \ D_{k}(\alpha,h)\mathcal{M}^{k}_{z\leftarrow z_{0}}(\alpha,h)=0,$ (5.54)
@@
$\begin{split}&\frac{1}{2\pi i}\operatorname{tr}_{L^{2}(\mathbb{R}/\mathbb{Z};\mathbb{C}^{2})}\oint_{\alpha}D_{k}(\zeta,h)^{-1}\partial_{\zeta}D_{k}(\zeta,h)d\zeta\\ &\ \ \ \ \ \frac{1}{2\pi i}\operatorname{tr}_{\mathbb{C}^{2}}\oint_{\alpha}(\mathcal{M}^{k}_{1\leftarrow 0}(\zeta,h)-I)^{-1}\partial_{\zeta}\mathcal{M}^{k}_{1\leftarrow 0}(\zeta,h)d\zeta.\end{split}$ (5.55)
A systematic derivation of this formula follows for setting up a Grushin problem (see [DyZw19, §C.1]) for $D_{k}(\alpha,h)$ following [SZ07, (3.12)]. To pose it we introduce
@@
$\begin{gathered}R_{-}:\mathbb{C}^{2}\to H^{2}(\mathbb{R}/\mathbb{Z}),\ \ \ (R_{-}u_{-})(x):=\chi^{\prime}_{-}\mathcal{M}^{k}_{x\leftarrow 0},\\ R_{+}:H^{2}(\mathbb{R}/\mathbb{Z})\to\mathbb{C}^{2},\ \ \ R_{+}u:=u(0).\end{gathered}$ (5.56)
@@
$\mathscr{P}^{-1}=\mathscr{E}:=\begin{pmatrix}E&E_{-}\\ E_{+}&E_{-+}\end{pmatrix},\ \ \ E_{-+}=I_{\mathbb{C}^{2}}-\mathcal{M}^{k}_{1\leftarrow 0}(\alpha,h).$ (5.58)
@@
To show invertibility of (5.56) it is enough to construct a left inverse. The operator $D_{k}(\alpha,h)$ is a Fredholm operator and as
@@
$\begin{split}\chi^{\prime}_{-}([x])I_{+}([x])v_{+}&=\chi^{\prime}_{-}(x)\mathcal{M}^{k}_{x\leftarrow 0}=\chi_{-}^{\prime}(x)\mathcal{M}^{k}_{x-1\leftarrow-1}=\chi_{-}^{\prime}(x)\mathcal{M}^{k}_{x-1\leftarrow 0}\mathcal{M}^{k}_{0\leftarrow-1}\\ &=\chi_{-}^{\prime}(x)I_{-}([x])\mathcal{M}^{k}_{1\leftarrow 0}.\end{split}$

### v2

@@ 5.1. Existence of protected states
Under an additional assumption it is easy to that there are no protected states at $k\in\Lambda^{*}$. That assumption, (5.1), can never hold for $V$ satisfyingthe symmetry assumptions (2.5).
@@
$\begin{gathered}\begin{pmatrix}(2D_{\bar{z}})^{2}&R_{-}\\ R_{+}&0\end{pmatrix}:H^{2}(\mathbb{C}/\Lambda;\mathbb{C})\oplus\mathbb{C}\to L^{2}(\mathbb{C}/\Lambda;\mathbb{C})\oplus\mathbb{C},\\ R_{-}u_{-}=u_{-},\ \ \ R_{+}u=\langle u,1\rangle.\end{gathered}$
@@
$\begin{pmatrix}E(\alpha)&E_{+}(\alpha)\\ E_{-}(\alpha)&E_{-+}(\alpha)\end{pmatrix},\ \ E_{-+}(\alpha)=\sum_{k=1}^{\infty}\alpha^{2k}E_{-}V(EV)^{k-1}E_{+}.$ (5.2)
@@
Remark. Suppose that $\Gamma\subset\Lambda^{*}$ such that $0\notin\Gamma$ and for $\gamma,\gamma^{\prime}\in\Gamma$, $\gamma+\gamma^{\prime}\in\Gamma$. Then consider
$V(z):=\sum_{\gamma\in\Gamma}a_{\gamma}e^{i\langle z,\gamma\rangle}\in C^{\infty}(\Lambda/\Gamma),\ \ \ a_{\gamma}=\mathcal{O}(|\gamma|^{-\infty}).$ (5.3)
@@
5.2. Basic properties of the toy model
We first assume that the potential takes the form $V(z)=W(z)^{2}$, where $W$ is periodic with respect to some lattice $\Lambda\subset\mathbb{C}$. In that case the characteristic variety $q^{-1}(0)$, $q=(2\bar{\zeta})^{2}-W(z)^{2}$ is given by a union of two disjoint tori $\{(x,\xi):2\bar{\zeta}=\pm W(z)\}$, $\zeta=\frac{1}{2}(\xi_{1}-i\xi_{2})$. These tori are Lagrangian if and only of $\operatorname{Im}\partial_{z}W(z)\equiv 0$ but we are interested in the general case of complex trigonometric polynomials $W(z)$ satisfying the condition
Figure 14. The characteristic variety $q^{-1}(0)$ for the classical symbol of (1.1), $q(x,\xi)=(2\bar{\zeta})^{2}-V(z)$ with $V(z)=W(z)^{2}$ and $W(z)\neq 0$. It is a union of two disjoint tori projecting diffeomorphically to the base torus. The tori are Lagrangian only if $\operatorname{Im}\partial_{z}W(z)\equiv 0$. Flat bands do not occur for this model but we can ask for the values of $\alpha$ for which $\ker_{H^{2}(\mathbb{C}/\Lambda)}P(\alpha,k)\neq\{0\}$. An answer in a special case is given in Theorem 3.
In view of Lemma 9 it natural to ask about the set
@@ 5.2. Basic properties of the toy model
$m(\alpha,k):=\frac{1}{2\pi i}\operatorname{tr}_{L^{2}(\mathbb{C}/\Lambda)}\oint_{\alpha}P(\beta,k)^{-1}\partial_{\beta}P(\beta,k)d\beta,$ (5.6)
where the integral is over a positively oriented circle including $\alpha$ as the only possible pole of $\beta\mapsto P(\beta,k)^{-1}$. If $k\in\Lambda^{*}$ this assumes meromorphy of $\beta\mapsto P(\beta,k)^{-1}$.
We henceforth specialize to the simplest case $\Lambda=\mathbb{Z}^{2}$ and $W(x,y)=W(x)$, $(x,y)\in\mathbb{R}^{2}$ where $W(x)=W(x+1)$ is real analytic and in fact extends to an entire function. That amounts to the study of the operator $P_{\zeta}(\alpha)$ defined in (1.13). We present a self-contained argument which indicates that even in the simplified setting (compared to that in §4) a rigorous derivation becomes quite involved.
As explained in (5.50) below, we can reduce the study of the kernel of $P_{\zeta}(\alpha)$ on $1$-periodic functions on $\mathbb{R}$ to the semiclassically rescaled operator $P(\alpha,h):=(hD_{x})^{2}-\alpha^{2}W(x)^{2}$ on functions on $\mathbb{R}$ satisfying $u(x+1)=e^{ik/h}u(x)$.
We now assume that the potential $W(x)$ extends to an entire function $W(z)$, which still has to be 1-periodic. The operator $P(\alpha,h)$ naturally extends to an operator on entire functions. This is the setting in which the analysis of this section is carried out. We emphasize that the meaning of the letter $z$ is different here than in the other sections: it is the complexification of $x$. Accordingly, we are looking for elements of the kernel of $P(\alpha,h)$ which are holomorphic in $z$.
@@ 5.3. Semiclassical formulation
for some constant $C_{0}$. We should mention that the problem scales and we could consider $|\alpha|\gg 1$ instead of (5.7) and $h\to 0$. It is however natural in the context of WKB to use real $h\to 0$ as the asymptotic parameter.
@@
For $z_{0},z_{1}\in\mathbb{C}$, define the transition matrix $\mathcal{M}_{z_{1}\leftarrow z_{0}}:\mathbb{C}^{2}\to\mathbb{C}^{2}$ as follows: for each solution $u$ to (5.8) we have
@@ 5.4. WKB approximation
We will now study the solutions to (5.8) using the global phase function $\Phi$.
@@ 5.4.2. WKB solutions
We next construct approximate WKB solutions, $u^{\pm}_{N}$, to the equation (5.8). Let $\gamma$ be a curve satisfying (5.15). Since $\gamma$ does not self intersect, we can fix a simply connected open set set $\mathcal{U}_{\gamma}$ such that
@@
$\begin{split}\operatorname{Wr}(u_{N}^{+},u_{N}^{-})&=e^{b_{N}^{+}+b_{N}^{-}}\left(\partial_{z}(-\alpha\Phi-ihb_{N}^{-})-\partial_{z}(\alpha\Phi-ihb_{N}^{+})\right)\\ &=\alpha e^{b_{N}^{+}+b_{N}^{-}}(-2W+i\tilde{h}\partial_{z}(b_{N}^{+}-b_{N}^{-})).\end{split}$
On the other hand, (5.12), (5.22) and (5.23) give $\partial_{z}\operatorname{Wr}(u_{N}^{+},u_{N}^{-})=\mathcal{O}(\tilde{h}^{N+1})$.
@@ 5.4.3. Approximating the transition matrix
$S_{\gamma,N}(\alpha,h):=\begin{pmatrix}e^{i(\Phi^{+}_{N}(\gamma(0);\alpha,h)-\Phi^{+}_{N}(\gamma(1);\alpha,h))/h}&0\\ 0&e^{i(\Phi^{-}_{N}(\gamma(0);\alpha,h)-\Phi^{-}_{N}(\gamma(1);\alpha,h))/h}\end{pmatrix}$. (5.29)
@@
$\mathcal{M}_{\gamma(1)\leftarrow\gamma(0)}=$ $\,B_{N}(\gamma(1);\alpha,h)^{-1}S_{\gamma,N}(\alpha,h)B_{N}(\gamma(0);\alpha,h)$ (5.30)
@@
$e^{-i\Phi^{+}_{N}(\gamma(1);\alpha,h)/h}\operatorname{Wr}_{+}(1)$ $=e^{(-i\Phi^{+}_{N}(\gamma(1);\alpha,h)+i\Phi^{+}_{N}(\gamma(0);\alpha,h))/h)}e^{-i\Phi^{+}_{N}(\gamma(0);\alpha,h)}\operatorname{Wr}_{+}(t)$
@@
$e^{-i\Phi^{-}_{N}(\gamma(1);\alpha,h)/h}\operatorname{Wr}_{-}(1)$ $=e^{(-i\Phi^{-}_{N}(\gamma(1);\alpha,h)+i\Phi^{-}_{N}(\gamma(0);\alpha,h))/h)}e^{-i\Phi^{-}_{N}(\gamma(0);\alpha,h)}\operatorname{Wr}_{+}(t)$
@@
(we used here the facts that $\operatorname{Re}(-i\Phi^{\pm}_{N}(\gamma(1);\alpha,h))=\pm\psi_{\gamma,\alpha}(1)+\mathcal{O}(h)$). That is,
${\rm{\bf{Wr}}}_{N}(\gamma(1))=S_{\gamma,N}(\alpha,h){\rm{\bf{Wr}}}_{N}(\gamma(0))+\mathcal{O}_{\gamma,N}(h^{N+1}e^{(\psi_{\gamma,\alpha}(1)-\psi_{\gamma,\alpha}(0))/h}|\mathbf{u}(\gamma(0))|).$
@@
$B_{N}(\gamma(1);\alpha,h)\mathbf{u}(\gamma(1))=$ $S_{\gamma,N}(\alpha,h)B_{N}(\gamma(0);\alpha,h)\mathbf{u}(\gamma(0))$
@@
$\mathbf{u}(\gamma(1))=$ $B_{N}(\gamma(1);\alpha,h)^{-1}S_{\gamma,N}(\alpha,h)B_{N}(\gamma(0);\alpha,h)\mathbf{u}(\gamma(0))$
@@ 5.5. WKB analysis in the periodic case
We also assume that $\gamma$ is a non-self-intersecting curve as in (5.15) such that
@@ 5.5.2. Stokes loops
Assume that $W$ is a 1-periodic entire function on $\mathbb{C}$. Let $\gamma:[0,1]\to\mathbb{C}$ be a smooth non-self-intersecting curve. We say that $\gamma$ is a Stokes loop for $W$, if $\gamma$ extends to a smooth non-self-intersecting curve $\widetilde{\gamma}$$\mathbb{R}\to\mathbb{C}$ and the following conditions hold for all $t\in\mathbb{R}$:
@@
$W(\gamma(t))\gamma^{\prime}(t)$ $>0.$ (5.40)
@@
Suppose that $\gamma$ is a Stokes loop in the sense of Definition 15. Then the winding number condition (5.34) holds.
@@
$\gamma(t):=t+if(t),\ \ f\in C^{\infty}(\mathbb{R};\mathbb{R}),\quad f(t+1)=f(t).$ (5.46)
@@ 5.6. Mutliplicity in terms of the transition matrix
$\begin{gathered}P_{k}(\alpha,h)=P(\alpha,h):H^{2}_{k}\to L^{2}_{k},\ \ \ H^{s}_{k}:=\{w\in H^{s}_{\rm{loc}}(\mathbb{R}):w(x+1)=e^{ik/h}w(x)\},\end{gathered}$ (5.48)
($H_{k}^{0}=L^{2}_{k}\simeq L^{2}(\mathbb{R}/\mathbb{Z})$) and then put
$m_{h}(\alpha,k)=\frac{1}{2\pi i}\operatorname{tr}_{L^{2}_{k}}\oint_{\alpha}P_{k}(\beta,h)^{-1}\partial_{\beta}P_{k}(\beta,h)d\beta,$ (5.49)
where the integral is over a positively oriented circle including $\alpha$ as the only possible pole of $\beta\mapsto P_{k}(\beta,h)$.
We note that, in the notation of (1.13),
$\begin{gathered}P_{k}(\alpha,h)=h^{2}\tau(k)P_{k/h}(\alpha/h)\tau(k)^{-1},\\ \tau(k)w(x):=e^{ikx/h}w(x),\ \ \ \tau(k):H^{s}(\mathbb{R}/\mathbb{Z})\to H^{s}_{k}.\end{gathered}$ (5.50)
Hence, the meromorphy of $\beta\mapsto P_{k}(\beta,h)$ is guaranteed by Lemma 9, except for $k/h\in 2\pi\mathbb{Z}$ when we can have $\ker P_{k}(\alpha,h)\neq\{0\}$ for all $\alpha$ (see the remark after Lemma 9). In that case, we put $m_{h}(\alpha,k)=\infty$.
Since we consider $V(x,y)=W(x)^{2}$, the Fourier mode decomposition in $y$ gives, in the notation of (1.1),
$P(\beta,k)=\bigoplus_{m\in\mathbb{Z}}P_{2\pi mi+k}(\beta).$
From (5.50) we then have
$m(\alpha/h,k/h)=\sum_{m\in\mathbb{Z}}m_{h}(\alpha,2\pi imh+k),$ (5.51)
@@
$m_{h}(\alpha,k)=\frac{1}{2\pi i}\operatorname{tr}_{\mathbb{C}^{2}}\oint_{\alpha}(\mathcal{M}_{z_{0}+1\leftarrow z_{0}}(\beta,h)-e^{ik/h})^{-1}\partial_{\beta}\mathcal{M}_{z_{0}+1\leftarrow z_{0}}(\beta,h)d\beta,$ (5.52)
@@
$\beta\mapsto\det(\mathcal{M}_{z_{0}+1\leftarrow z_{0}}(\beta,h)-e^{ik/h})$
at $\beta=\alpha$.
@@
1. We start by observing that since $W(z+1)=W(z)$, $\mathcal{M}_{z_{0}\leftarrow z}=\mathcal{M}_{z_{0}+1\leftarrow z+1}$. Hence
@@
and the statements for a fixed $z_{0}$ are equivalent to statements for all $z$.
The obstruction to the meromorphy of $P_{k}(\alpha,h)^{-1}$ is having a non-trivial solution $w(\alpha)\in H_{k}^{2}$ to $P_{k}(\alpha,h)w(\alpha)=0$ for all $\alpha$. As discussed in §5.2, $x\mapsto w(\alpha,x)$ extends to an entire function which then satisfies (5.47) and $P(\alpha,h)w=0$. Definition of $\mathcal{M}_{z+1\leftarrow z}(\alpha,h)$ in §5.3, then shows that $e^{ik/h}$ is an eigenvalue of $\mathcal{M}_{z+1\leftarrow z}$ for all $z$. That gives $\Leftarrow$ in the first statement. To see $\Rightarrow$ we note that meromorphy of $P_{k}(\alpha,h)^{-1}$ implies existence of $\alpha_{0}$ for which $\ker P_{k}(\alpha_{0},h)=\{0\}$. If $e^{ik/h}$ were an eigenvalue of $\mathcal{M}_{z_{0}+1\leftarrow z_{0}}(\alpha_{0},h)$ then, by (5.53) it would be an eigenvalue of $\mathcal{M}_{z+1\leftarrow z}(\alpha_{0},h)$ for all $z$ and we would have have an entire solution to $P(\alpha_{0},h)u=0$ satisfying (5.47). But that means that $\ker P_{k}(\alpha,h)$ is nontrivial.
@@
The related operator acting on holomorphic functions is given by $\widetilde{P}_{k}(\alpha,h)=(hD_{z}+k)^{2}-\alpha^{2}W^{2}$. The corresponding transition matrix, $\mathcal{M}^{k}_{z_{1}\leftarrow z_{0}}$, is defined as in (5.10) but with (5.9) replaced by $\mathbf{u}(z)=(u(z),(hD_{z}+k)u(z))$. It is related to the original transition matrix as follows:
$\mathcal{M}^{k}_{z_{1}\leftarrow z_{0}}(\alpha,h)=e^{ik(z_{0}-z_{1})/h}\mathcal{M}_{z_{1}\leftarrow z_{0}}(\alpha,h).$ (5.54)
@@
$D_{k}(\alpha,h):=\begin{pmatrix}hD_{z}+k&-1\\ -\alpha^{2}W(z)^{2}&hD_{z}+k\end{pmatrix},\ \ \ D_{k}(\alpha,h)\mathcal{M}^{k}_{z\leftarrow z_{0}}(\alpha,h)=0,$ (5.55)
@@
$\begin{split}&\frac{1}{2\pi i}\operatorname{tr}_{L^{2}(\mathbb{R}/\mathbb{Z};\mathbb{C}^{2})}\oint_{\alpha}D_{k}(\beta,h)^{-1}\partial_{\beta}D_{k}(\beta,h)d\beta=\\ &\ \ \ \ \ \frac{1}{2\pi i}\operatorname{tr}_{\mathbb{C}^{2}}\oint_{\alpha}(\mathcal{M}^{k}_{1\leftarrow 0}(\beta,h)-I)^{-1}\partial_{\beta}\mathcal{M}^{k}_{1\leftarrow 0}(\beta,h)d\beta.\end{split}$ (5.56)
2. A systematic derivation of the formula (5.56) follows for setting up a Grushin problem (see [DyZw19, §C.1]) for $D_{k}(\alpha,h)$ following [SZ07, (3.12)]. To pose it we introduce
@@
$\begin{gathered}R_{-}:\mathbb{C}^{2}\to L^{2}(\mathbb{R}/\mathbb{Z};\mathbb{C}^{2}),\ \ \ (R_{-}u_{-})(x):=\chi^{\prime}_{-}\mathcal{M}^{k}_{x\leftarrow 0}u_{-},\\ R_{+}:H^{1}(\mathbb{R}/\mathbb{Z};\mathbb{C}^{2})\to\mathbb{C}^{2},\ \ \ R_{+}\mathbf{u}:=\mathbf{u}(0).\end{gathered}$ (5.57)
@@
$\mathscr{P}^{-1}=\mathscr{E}:=\begin{pmatrix}E&E_{-}\\ E_{+}&E_{-+}\end{pmatrix},\ \ \ E_{-+}=\mathcal{M}^{k}_{0\leftarrow 1}-I_{\mathbb{C}^{2}}(\alpha,h).$ (5.59)
@@
3. To show invertibility of (5.57) it is enough to construct an inverse. The operator $D_{k}(\alpha,h)$ is a Fredholm operator and as
@@
$\begin{split}\chi^{\prime}_{-}([x])I_{+}([x])&=\chi^{\prime}_{-}(x)\mathcal{M}^{k}_{x\leftarrow 0}=\chi_{-}^{\prime}(x)\mathcal{M}^{k}_{x-1\leftarrow-1}=\chi_{-}^{\prime}(x)\mathcal{M}^{k}_{x-1\leftarrow 0}\mathcal{M}^{k}_{0\leftarrow-1}\\ &=\chi_{-}^{\prime}(x)I_{-}([x])\mathcal{M}^{k}_{1\leftarrow 0}.\end{split}$